I am crashing and burning in my college algebra class...
Write the equation of a line that passes (2,-5) and (3,-10). Write the answer in slope-intercept form.
I have determined that the slope is -15 and I know the slope-intercept formula is y = mx + b....
what do I do next???

Answer by queenofit(120)

If you have determined the slope of the line
M= Y2-Y1
-------- = -10 - (-5) / 3-2 = -15
X2-X1
Then use this formula first
Y-Y1= m(X-X1)
y- (-10)= -15(x-2)
y + 10 = -15x + 30
Y= -15X +30-10
Y= -15X +20 (SLOPE INTERCEPT FORM).